# DynaSpec: Context-aware Dynamic Speculative Sampling for Large-Vocabulary Language Models

### v1

@@ 4.2 Meta classifiers: Context-aware token Selection
0: prefix sequence $c$, speculation length $\gamma$, target model $\mathcal{T}$, router $r_{\theta}$, drafter $\mathcal{D}$, LM head $W_{LM}$, clusters $\{C_{m}\}$ over vocabulary, first token cluster budget $k_{\mathrm{max}}$, token budget $k_{\mathrm{t}}$, empty draft list $d$, empty draft list score d_scoresd\_scores
@@
3:   (stream $S_{m}$) $s\leftarrow r_{\theta}([h_{c[-1]},emb(x_{j})])$; $\mathcal{K}\leftarrow\operatorname{TopK}_{k_{\mathrm{max}}}(s)$; $I\leftarrow\mathrm{indices}\!\big(\cup_{m\in\mathcal{K}}C_{m}\big)$
4:   (stream $S_{d}$) $\tilde{h}\leftarrow\mathcal{D}\!\big([h_{c},c[1:]+x_{j}]\big)$
5:   $last\_step\_scores\leftarrow 0$
6:  else
7:   if $j>1$ then $k_{\mathrm{c}}\leftarrow k_{\mathrm{max}}//((i+1)*2)$ else $k_{\mathrm{c}}\leftarrow k_{\mathrm{max}}$
8:   (stream $S_{m}$) $s\leftarrow r_{\theta}([h_{j-1},emb(x_{j})])$; $\mathcal{K}\leftarrow\operatorname{TopK}_{k_{\mathrm{c}}}(s)$; $I\leftarrow\mathrm{indices}\!\big(\cup_{m\in\mathcal{K}}C_{m}\big)$
9:   (stream $S_{d}$) $\tilde{h}\leftarrow\mathcal{D}\!\big([h_{j-1},x_{j}]\big)$
10:  end if
@@
11:  sync $S_{m},S_{d}$; $z\leftarrow\mathrm{FUSED\_INDEX\_GEMM}\!\big(\tilde{h},\ W_{LM},\ I\big)$ // shortlist logits on gathered columns II
@@
12:  $p\leftarrow\operatorname{log\_softmax}(z)$;  $T_{j},TopP_{j}\leftarrow\operatorname{TopK}_{k_{\mathrm{t}}}(p)$;  $\tilde{T_{j}}=remap2realid(T_{j})$
13:  $cu\_scores\leftarrow TopP_{j}+last\_step\_scores$
@@
14:  $d\leftarrow d\oplus\tilde{T}_{j}$$d\_score\leftarrow d\_score\oplus cu\_scores$ // append score and token to the draft list
@@
15:  $TopC_{j},last\_step\_scores\leftarrow TopK_{k_{\mathrm{t}}}(cu\_scores)$
16:  $x_{j}\leftarrow\tilde{T}_{j}[TopC_{j}]$
17:  if $j=0$ then $h_{j}\leftarrow\tilde{h}[-1]$ else $h_{j}\leftarrow\tilde{h}[TopC_{j}]$
18:end for
19: Re-ranking the $d$ based on their corresponding scores d_scoresd\_scores
20: verify once with $\mathcal{T}$ over full $\mathcal{V}$ on the top-$k$ subset of $d$;
4.3 Dynamic Drafting
At each step $t$, the router (meta-classifier) selects the shortlist indices, while the drafter generates the hidden state for that step. A single gathered-head GEMM is then performed over the shortlisted tokens using this hidden state. The router runs on a separate CUDA stream and typically finishes before the drafter block. The total draft time for a single pass can be expressed as,
$T_{D}(c,t)\ \approx\ T_{\text{embed}}+\max\!\big\{T_{\text{core}},\,T_{\text{meta}}\big\}\ +T_{\text{index+gemm}}\!\big(B(c,t)\big),$
where $T_{\text{meta}}\!\ll\!T_{\text{core}}$ and $T_{\text{embed}}$ is generally small. Thus the only vocabulary-dependent term is the $d\times B(c,t)$ head compute (with $B(c,t)\!\ll\!|\mathcal{V}|$), reducing the denominator in Eq. 2 while preserving acceptance via context-aware support. Since $B(c,t)\!\propto\!|\mathcal{V}_{S}(c,t)|$ and the position-aware budget $k_{\mathrm{c}}(t)$ decay with $t$, the shortlist and thus the head compute shrink as generation proceeds, while larger early budgets preserve acceptance. In our implementation, index selection and the head GEMM are fused into a single custom CUDA kernel to avoid intermediate buffers and redundant global-memory traffic: the kernel performs warp-level column gathers with asynchronous, double-buffered shared-memory tiling, computes on Tensor Cores, and uses a persistent L2 access-policy window to keep activations cached. Verification is unchanged from standard speculative decoding.
5 Experimental Results
5.1 Experimental Settings
We adhere to the experimental setup and evaluation of the EAGLE (Li et al., 2024a, b) family and FR-Spec (Zhao et al., 2025). We implement the baseline FR-Spec and DynaSpec in the EAGLE 2 framework.
Datasets and Models. We benchmark DynaSpec over a diverse mix of seven tasks. Spec-Bench (Xia et al., 2024) supplies six of them—Machine Translation (WMT14 DE to EN) (Bojar et al., 2014), Multi-turn Conversation (MT-Bench) (Zheng et al., 2023), RAG and QA from Natural Questions (Kwiatkowski et al., 2019), Math from GSM8K (Cobbe et al., 2021), and Summarization (CNN/DailyMail) (Nallapati et al., 2016)—each evaluated on 80 prompts. The seventh task is Code, for which we employ HumanEval (Chen et al., 2021) with 164 problems. We bound generations to $1024$ tokens for Spec-Bench subtasks and HumanEval. We used Llama series of models, specifically Llama-33-8B-Instruct, which uses a vocabulary size of $128K$ for our experiments.
Evaluation metrics and Hardware Settings. We evaluate using metrics: #Mean Acceptance Length, the average number of tokens committed per draft–verification cycle. #Mean Acceptance Length is hardware-agnostic but does not capture the overhead of drafting. Our primary baseline is FR-Spec (Zhao et al., 2025), a static subset–selection approach built on EAGLE-2 (Li et al., 2024b); in contrast, DynaSpec performs dynamic subset selection atop EAGLE-2 (Li et al., 2024b). All experiments were run on a single NVIDIA A6000 GPU.
@@ 5.2 Mean Acceptance Length
Table 1 presents a comparison between DynaSpec and FR-Spec. We observe that DynaSpec achieves a higher mean acceptance length than FR-Spec while requiring a smaller mean shortlist size across seven datasets. Consistent with the findings of EAGLE-2, we also note that QA and summarization tasks yield lower accepted lengths compared to other tasks.
Config MT Conv. RAG Math QA Summ. Code Avg. Full Vocab $3.66$ $4.11$ $4.03$ $4.31$ $3.45$ $3.68$ $4.77$ $4.00$ vocab size=128k FR-Spec $3.38$ $3.87$ $3.85$ $4.16$ $3.32$ $3.51$ $4.11$ $3.74$ vocab size=32k DynaSpec $3.51$ $4.05$ $3.91$ $4.21$ $3.40$ $3.51$ $4.71$ $3.90$ vocab size 27,817 27,020 27,546 27,436 27,208 27,107 27,277 27,344
Table 1: Mean accepted tokens for Llama-3-8B on seven datasets. Vocab size indicates the shortlisted vocabulary size. Full Vocab (128k) corresponds to the EAGLE-2 baseline; FR-Spec uses 32k. For Dyna-Spec, shortlist size varies by dataset; per-dataset means are shown individually.
@@ 5.3 Position-Aware Vocabulary Shortlisting
For the position-aware strategy, the results are presented in Figure 2. It shows that using a position-aware subset size outperforms a fixed size across all draft steps, as the first few tokens are more important than the later ones—an observation consistent with Gumiho Li et al. (2025). Table 2 further reports the results of combining the position-aware strategy with FR-Spec, which also demonstrates the effectiveness of this approach.
@@ 6 Conclusion
We address a central bottleneck in speculative decoding for large-vocabulary LLMs—the drafter’s $\mathcal{O}(|\mathcal{V}|d)$ output head. DynaSpec replaces static shortlists with a context-aware dynamic head that routes each prefix to a small number of token clusters and evaluates only over their union, while verification remains exact on the full vocabulary. Our analysis shows that context-conditioned supports strictly dominate any fixed subset in expected acceptance, and BCE-with-top-$k$ routing maximizes a lower bound on retained target mass. Systems-wise, we fuse index selection with the head GEMM and run routing in parallel, reducing draft-time cost. Empirically, DynaSpec increases mean acceptance length over fixed-shortlist baselines across tasks and models, without altering the target distribution. The method is plug-in compatible with EAGLE-style pipelines and orthogonal to cache reuse and serving-stack optimizations.
@@ A.2 Meta classifier training
To train the meta classifier, we run eagle-2 Li et al. [2024b] on subsets of ShareGPT and UltraChat200K Ding et al. [2023], and collect the hidden states of the draft model, token embeddings, and the top-k tokens at each draft step. Using these, we construct the training dataset for the meta-classifier as $\{(h_{i-1},embed(x_{i})),map2meta(top\_k\_tokens_{i})\}$, where $h_{i}$ is the hidden state of the draft model at step $i$, $embed(x_{i})$ is the embedding of token $x_{i}$, and $map2meta(top\_k\_tokens_{i})$ represents the cluster IDs obtained by mapping the top-k tokens to their clusters.

### v2

@@ 5.1 Experimental Settings
0: prefix sequence $c$, speculation length $\gamma$, target model $\mathcal{T}$, router $r_{\theta}$, drafter $\mathcal{D}$, LM head $W_{LM}$, clusters $\{C_{m}\}$ over vocabulary, first token cluster budget $k_{\mathrm{first}}$, second token cluster budget $k_{\mathrm{second}}$, token budget $k_{\mathrm{t}}$, empty draft list $d$, empty draft list score $\mathrm{d\_scores}$
@@
3:   (stream $S_{m}$) $s\leftarrow r_{\theta}([h_{c[-1]},emb(x_{j})])$;
4:   $\mathcal{K}\leftarrow\operatorname{TopK}_{k_{\mathrm{first}}}(s)$; $I\leftarrow\mathrm{indices}\!\big(\cup_{m\in\mathcal{K}}C_{m}\big)$
5:   (stream $S_{d}$) $\tilde{h}\leftarrow\mathcal{D}\!\big([h_{c},c[1:]+x_{j}]\big)$
6:   $\mathrm{last\_step\_scores}\leftarrow 0$
7:  else
8:   if $j>1$ then $k_{\mathrm{c}}\leftarrow k_{\mathrm{second}}//(i*2)$ else $k_{\mathrm{c}}\leftarrow k_{\mathrm{second}}$
@@
9:   (stream $S_{m}$) $s\leftarrow r_{\theta}([h_{j-1},\mathrm{emb}(x_{j})])$; $\mathcal{K}\leftarrow\operatorname{TopK}_{k_{\mathrm{c}}}(s)$; $I\leftarrow\mathrm{indices}\!\big(\cup_{m\in\mathcal{K}}C_{m}\big)$
10:   (stream $S_{d}$) $\tilde{h}\leftarrow\mathcal{D}\!\big([h_{j-1},x_{j}]\big)$
11:  end if
12:  sync $S_{m},S_{d}$;
@@
13:  $z\leftarrow\mathrm{FUSED\_INDEX\_GEMM}\!\big(\tilde{h},\ W_{LM},\ I\big)$ // shortlist logits on gathered columns II
@@
14:  $p\leftarrow\operatorname{log\_softmax}(z)$;  $T_{j},TopP_{j}\leftarrow\operatorname{TopK}_{k_{\mathrm{t}}}(p)$;  $\tilde{T_{j}}=\operatorname{remap2realid}(T_{j})$
15:  $\mathrm{cu\_scores}\leftarrow\operatorname{TopP}_{j}+\mathrm{last\_step\_scores}$
@@
16:  $d\leftarrow d\oplus\tilde{T}_{j}$$\mathrm{d\_score}\leftarrow\mathrm{d\_score}\oplus\mathrm{cu\_scores}$ // append score and token to the draft list
@@
17:  $\operatorname{TopC}_{j},\mathrm{last\_step\_scores}\leftarrow\operatorname{TopK}_{k_{\mathrm{t}}}(\mathrm{cu\_scores})$
18:  $x_{j}\leftarrow\tilde{T}_{j}[\operatorname{TopC}_{j}]$
19:  if $j=0$ then $h_{j}\leftarrow\tilde{h}[-1]$ else $h_{j}\leftarrow\tilde{h}[\operatorname{TopC}_{j}]$
20:end for
21: Re-ranking the $d$ based on their corresponding scores $\mathrm{d\_scores}$
22: verify once with $\mathcal{T}$ over full $\mathcal{V}$ on the top-$k$ subset of $d$;
@@ 5.2 Mean Acceptance Length
Table 1 summarizes mean acceptance lengths across seven benchmarks using Llama-3-8B. DynaSpec improves average acceptance from $3.64$ (FR-Spec) to $3.79$ tokens per verification step, about a 4.1% relative gain, while using a smaller average shortlist of roughly 27K tokens compared to FR-Spec’s fixed 32K. This indicates that context-aware routing recovers most of the full-vocabulary performance ($3.89$) at substantially lower draft-time cost. Consistent with observations from EAGLE-2, tasks such as QA and summarization exhibit shorter accepted lengths. In contrast, code generation, where rare tokens are crucial, shows the largest gap: FR-Spec drops to $3.31$, while DynaSpec achieves $3.85$, nearly matching full-vocabulary decoding ($3.92$). These results highlight that dynamic, context-dependent shortlisting preserves acceptance quality even under vocabulary reduction. Similar trends hold for Qwen-2-7B (Table 1), confirming the robustness of the method across architectures.
5.3 Decoding Speed
We evaluate end-to-end decoding throughput using the CUDA-based implementation of FR-Spec and our plug-in meta-classifier extension. Results are summarized in Table 2. DynaSpec achieves higher token-generation throughput than FR-Spec on four of the seven datasets, confirming that dynamic, context-aware shortlisting effectively reduces draft-time overhead. For the remaining tasks, slightly lower speeds are observed, which we attribute to the current implementation of the sub-sampled (indexed) matrix multiplication kernel. Further kernel-level optimization is expected to close this gap. Overall, DynaSpec demonstrates consistent efficiency gains across diverse workloads while maintaining full-vocabulary exactness during verification.
@@ 5.4 Position-Aware Vocabulary Shortlisting
For the position-aware strategy, the results are presented in Figure 3. Using a position-aware subset size outperforms a fixed size across all draft steps, as the first few tokens have a disproportionate impact on acceptance and throughput, an observation consistent with prior findings in Gumiho li2025gumiho. Allocating a larger cluster budget to early tokens preserves acceptance where it matters most, while decaying the shortlist later reduces indexed-GEMM cost without affecting verification. As shown in Figure 3, this adaptive schedule yields higher mean accepted lengths at comparable or smaller shortlist sizes. Table 3 further reports results for combining the position-aware schedule with FR-Spec, which likewise improves mean acceptance and reduces average shortlist size. Overall, these results confirm that position awareness complements context-aware routing by focusing computation on the most influential draft steps and amortizing overhead in later ones.
@@ 6 Conclusion
[…]
@@ A.2 Meta classifier training
To train the meta classifier, we run eagle-2 li2024eagle2 on subsets of ShareGPT and UltraChat200K ding2023enhancing, and collect the hidden states of the draft model, token embeddings, and the top-k tokens at each draft step. Using these, we construct the training dataset for the meta-classifier as $\{(h_{i-1},embed(x_{i})),map2meta(top\_k\_tokens_{i})\}$, where $h_{i}$ is the hidden state of the draft model at step $i$, $embed(x_{i})$ is the embedding of token $x_{i}$, and $map2meta(top\_k\_tokens_{i})$ represents the cluster IDs obtained by mapping the top-k tokens to their clusters.